OpenStudy (anonymous):

what is bigger e^x or x^2

OpenStudy (nincompoop):

neither

OpenStudy (anonymous):

depends what x is

OpenStudy (anonymous):

If x=1:\[e^1>1^2\] If x=2:\[e^2>2^2\] If x=3:\[e^3>3^2\] Yet if x=-1: \[e^-1<(-1)^2\] Truth is, \[e^x>x^2\]If x is positive or 0.\[e^x<x^2\]If x is negative.

OpenStudy (anonymous):

\[ \lim_{x \rightarrow infinity } \frac{ f(x) }{ g(x) }=0\]

OpenStudy (anonymous):

The blue function is e^x, and the purple function is x^2. The blue is over the purple after 0, before 0 it is under.

OpenStudy (anonymous):

f(x)=e^x g(x)=2^x

OpenStudy (jack1):

@sarcos11 2nd equation is x^2 man... not 2^x hey